OpenStudy (anonymous):

3+sqrt{z-1} =sqrt{z+14} 3+sqrt{z-1} =sqrt{z+14} @Mathematics

OpenStudy (anonymous):

\[3+\sqrt{z-1 }= \sqrt{z+14}\]

OpenStudy (agreene):

z=2 but the way I got there.... I don't think you would know...

OpenStudy (anonymous):

\[3+\sqrt{z-11} = \sqrt{z+4}\]

OpenStudy (anonymous):

i got 2.

OpenStudy (anonymous):

i have to square both terms then use foil

OpenStudy (anonymous):

im just making sure they are correct cause i always make mistakes

OpenStudy (anonymous):

i have no clue how i posted the other problem

OpenStudy (agreene):

\[3+\sqrt{z-11} = \sqrt{z+4}\] \[z=12\] I've been using using phasor form and de moivre's

OpenStudy (anonymous):

oh yea i dont know what that is lol

OpenStudy (agreene):

basically, it's a quicker route then trying to get everything to FOIL correctly... but it involves switching to sin and cos ... which can be a bit more difficult if you're not used to it.

OpenStudy (anonymous):

oh yea i would totally get lost lol i have a hard time doing this darn math